### v1

@@ 2.1 Blocks
Ethereum subdivides time into 12-second slots, during which one block at most can be added to the blockchain. For our purposes, each slot corresponds to a bidding cycle. Slot numbers differ from block numbers; for example, our dataset covers all blocks submitted via the Agnostic relay from slot 10,534,387 to 10,534,425, corresponding to winning blocks 21,322,622 to 21,322,660. For ease of exposition, in what follows, we use winning block numbers whenever possible. Also relevant is that a relay will not hold an auction for a given slot if the proposer assigned to that slot is not registered with the relay. This situation appears in our data: out of the 39 slots we consider, we observe no bids for 11 bidding cycles because the proposers were not connected with the Agnostic relay. [^9] Specifically, we observe no bids for the auction cycles that led to the inclusion of blocks 21,322,624; 21,322,625; 21,322,627; 21,322,629; 21,322,633; 21,322,634; 21,322,636; 21,322,644; 21,322,646; 21,322,647; 21,322,651.
@@
However, when examining the proportion of blocks submitted by the top builders across all relays, we note an imbalance in our primary dataset. Across all relays, Titan Builder, submitted 13,679 blocks (25.6%), Rsync Builder 7,936 blocks (14.9%), Flashbots 2,626 blocks (4.9%), and Beaverbuild 21,447 blocks (40.2%). Hence, Beaverbuild submits proportionally fewer blocks to Agnostic Relay than to other relays. Our primary dataset is, therefore, an unbalanced subsample of all blocks submitted across all relays. We acknowledge this as a limitation of our data.
To explore why builders submit fewer blocks to Agnostic Relay than to other relays, we examine the timestamps associated with the blocks that were shared between Agnostic and at least one other relay. We find that only 5% were received by Agnostic before any other relay. More specifically, among the blocks received by both Agnostic and Ultrasound, just 1% of blocks were received by Agnostic first; between blocks received by both Agnostic and BloXroute only 15% were received by Agnostic first. Agnostic, therefore, has a latency disadvantage relative to the other major relays, which may explain why it receives fewer blocks from builders.
@@ 2.2 Transactions
A notable feature of our primary dataset is that 2,981 transactions — 35% of all unique transactions — originate from just two addresses. The first address,1111110x68d3A973E7272EB388022a5C6518d9b2a2e66fBf. is responsible for 1,873 transactions, of which only 31 were included in a winning block. Notably, only 6 (out of 1,873) transactions from this address appear in blocks not built by Titan, and all 6 are simple token transfers to Binance. The second address,1212120x51C72848c68a965f66FA7a88855F9f7784502a7F submitted 1,108 transactions, with just 97 included in a winning block. Only 249 (out of 1,108) transactions appear in blocks not built by Rsync, and only 4 appear in blocks not built by either Rsync or Titan. These 4 transactions are also simple token transfers, including two to addresses associated with Wintermute, an algorithmic trading fund that operates Rsync Builder. Based on these patterns, we interpret these addresses as searchers integrated with the Titan and Rsync builders, respectively. We refer to them as “Titan-bot” and “Rsync-bot”. [^13] Using a different methodology, Heimbach et al. (2024) also identified Rsync-bot as a non-atomic arbitrage bot, while Titan-bot did not yet exist during the period covered by their study.
@@
To further support our identification of “Rsync-bot” and “Titan-bot,” we analyze whether these searchers submit different transactions to different builders. Using transaction hashes, we find no case in which Rsync-bot sent the same transaction to both Rsync and Titan. However, when comparing transaction logs — i.e., the sequence of operations executed by the transaction and their outputs — we find 203 such cases. The discrepancy between transaction hashes and logs arises because Rsync-bot modifies the fee offered depending on the builder: on average, it pays 18% more to Rsync than to Titan for executing the same transaction. Moreover, we find that transactions sent exclusively to Rsync often include a provision that renders them non-executable if the block’s fee recipient is not Rsync. This behavior further supports our interpretation that Rsync-bot is integrated with Rsync and that the 18% fee differential represents the bot’s targeted profit margin when interacting with non-integrated builders. To perform the same analysis for Titan-bot, we consider winning blocks for our study period that are not in our primary dataset (because they were not submitted via Agnostic relay). Doing so, we find 20 transactions originating from Titan-bot included in blocks won by Beaverbuild. In 3 cases, the same transactions also appear in our dataset in blocks built by Titan. When comparing transaction logs, we find another two transactions in blocks built by Beaverbuild and Titan. In both cases, Titan-bot paid more for inclusion (by 43% and 23%), when the transaction was sent to Titan than Beaverbuild, again, supporting our interpretation that Titan-bot is integrated with Titan.
Almost all transactions from these bots involve token swaps on decentralized exchanges (DEXs): 1,864 for Titan-bot and 1,095 for Rsync-bot. For Titan-bot, the most frequently traded asset pairs are USDC/WETH (334 swaps), USDT/WETH (284 swaps), LINK/WETH (219 swaps), DAI/WETH (202 swaps), and MOGCOIN/WETH (200 swaps). The most-frequently traded asset pairs for Rsync-bot are instead USDT/WETH (315 swaps), USDC/WETH (208 swaps), WBTC/cbBTC (150 swaps), WBTC/WETH (81 swaps), and WBTC/USDT (55 swaps). We later discuss how these bots often trade on the same DEX pools during the same auction cycle. Based on this activity, we infer that both bots likely engage in non-atomic arbitrage — that is, arbitrage between on-chain and off-chain financial markets — and are often in competition.
@@
Swaps on decentralized exchanges (DEXes) are also the most common type of transaction across all unique transactions in our primary dataset. We identify 5,291 swap transactions, of which 3,919 occur on Uniswap V3, 1,228 on Uniswap V2, 142 on Sushiswap, and 101 on Pancakeswap. [^14] We also observe 99 transactions that interact with multiple DEXes, which is why the total number of swaps on individual DEXes exceeds the number of unique swap transactions. Despite their prevalence, only 814 swap transactions — approximately 15% of all swap transactions — were ultimately included on-chain. A small number of addresses, all belonging to searchers, account for the majority of swaps: the top four addresses are responsible for 3,653 swap transactions (about 70%), with Rsync-bot and Titan-bot alone responsible for 2,959 (56%). Swaps are also concentrated among a few pools: the top 11 DEX pools account for 2,872 transactions (54.4%). The most frequently traded token pairs are USDC/WETH, USDT/WETH, MOGCOIN/WETH, and LINK/WETH.
@@
We also identify 807 swap transactions that interact with 42 distinct DEX routers, services that optimize swap execution by splitting trades across multiple liquidity pools and/or using intermediate assets. The most frequently used routers are the Uniswap Universal Router (190 transactions), Uniswap V3 Router (172), 1inch V5 Router (52), and Uniswap V2 Router (42). Of these, 594 transactions were included in a winning block, implying that the majority of swaps included on-chain were routed through a DEX router. The remaining 213 router-based transactions were not included, and the vast majority of these (153) originate from a single user address, 0xf5213a6a2f0890321712520b8048d9886c1a9900. This user submitted 153 transactions calling the Uniswap V3 Router, of which only two were ultimately included in a winning block.
2.3 Transaction delivery
A notable feature of our dataset is that it allows us to distinguish between private transactions shared among multiple builders and exclusive transactions — that is, transactions available to only a single builder. As already discussed, this distinction is important in light of the current concentration in the builders’ market, which is often attributed to builders’ exclusive transactions.
We use data from MempoolGuru to label transactions as either public (i.e., they appeared in the public mempool) or non-public (i.e., they appeared in a block without first appearing in the public mempool). We also incorporate data from MEV Blocker and Flashbots Protect, two private mempools that share transactions with multiple builders (it is important to note that these public sources only classify transactions that were ultimately included in winning blocks).
We find that, of the 5,576 transactions in our primary dataset that were included in a winning block, MempoolGuru classifies 4,124 as public and 1,452 as non-public. Among the non-public transactions, we find that 266 were submitted through MEV Blocker, 40 through Flashbots Protect, and 14 through both — implying that these transactions are private but not exclusive. Using our dataset, we further investigate the remaining 1,132 non-public transactions included on-chain. We find that 642 of them — 45% of all non-public transactions in winning blocks — appear only in blocks proposed by a single builder, indicating that they are exclusive. Of these, 533 appear only in blocks by Titan Builder, 107 only in blocks by Rsync Builder, and 2 only in blocks by builder0x69. In terms of value, among the winning blocks, the sum of all fees paid to block builders for inclusion is 14.67 ETH, of which 12.45 ETH (85%) by exclusive transactions, 1.47 ETH (10%) by private transactions, and 0.75 ETH (5%) by public transactions. Exclusive transactions, therefore, are builders’ dominant source of revenues.
With respect to the 5,217 transactions never included in a winning block, about half (2,853 or 54.7%) are from Titan-bot and Rsync-bot, which, as we already discussed, are exclusive transactions (although, at times, Rsync-bot also sent transactions to Titan). Of the remaining 2,002 transactions, 1,554 (77.6%) transactions are present only in blocks by a single builder and are, therefore, exclusive. Hence, approximately 85% of the non-public transactions that were never included in a winning block are exclusive, roughly double the proportion of exclusive transactions among non-public transactions included in winning blocks.
As already discussed, one limitation of our analysis is that we observe only a small subset of blocks produced by Beaverbuild. As a result, some transactions we classify as “exclusive” may be present in some Beaverbuild blocks that are not included in our dataset. However, in Section 4 we analyze bilateral sharing of transactions between builders and find that while some sharing does occur, it involves relatively few transactions. If we extrapolate this result to transactions shared with Beaverbuild, our estimates of exclusivity should remain approximately accurate. A more important concern is that transactions exclusive to Beaverbuild do not appear in our primary dataset if Beaverbuild did not submit the corresponding block to Agnostic, regardless of whether the transaction was ultimately included in a winning block. This suggests that the share of exclusive transactions among all transactions submitted for inclusion during the study period may be larger than the share of exclusive transactions in our primary dataset.
@@ 3 Block 21322649
We now turn to a detailed empirical analysis of a specific auction cycle—namely, the one that resulted in the inclusion of block 21,322,649 on the Ethereum blockchain. This cycle is noteworthy for two reasons. First, bid data across all relays reveals intense competition between Titan Builder and Rsync Builder, with Rsync ultimately winning the auction (see Figures 2(a) and 2(b)). Second, Rsync-bot and Titan-bot actively compete in the same DEX pools, and account for 94%–97% of the total revenue generated by blocks in which they are included. We can therefore closely examine the dynamics of the bidding cycle by analyzing the two bots’ behavior and their respective builders’ bidding strategies.
(a) Builders’ bids across all relays during the auction cycle for Block 21322649
(b) Number of Unique transactions, blocks and swaps vs ETH price change on Binance, during the auction cycle for Block 21322649
Figure 2:
@@ 3.1 Description of the data for the auction cycle of block 21,322,649
For this auction cycle, our primary dataset contains 669 blocks and 504 unique transactions, of which 300 (59%) are swaps. Only 22 swaps were included in winning block 21,322,649, 8 from users and 14 from searchers. An additional 8 swaps were included in winning block 21,322,650, 3 more in winning block 21,322,651 and 1 more in winning block 21,322,652. Notably, all the swaps included in these later blocks are from users. A total of 181 transactions are attributed to Titan-bot and 42 to Rsync-bot, all of which are DEX swaps. None of the Titan-bot transactions were included in a winning block. In contrast, 5 transactions from Rsync-bot were included, all in winning block 21,322,649.
@@
As the bidding cycle progresses, we observe a sharp increase in the number of swaps (see Figure 2(b)). This is accompanied by a rise in volume swapped in each block (see Figure 3(a)), as well as in the share of swap volume attributable to the two main searchers, Titan-bot and Rsync-bot (see Figure 3(b)). Together, these two bots account for 94%–97% of the total revenues of blocks in which they are present, indicating that their activity is the primary driver of auction dynamics during this cycle. In contrast to public transactions, which arrive at a relatively uniform rate throughout the bidding cycle, private and, especially, exclusive transactions tend to arrive toward the end of the cycle (see Figure 3(c)). This pattern is also reflected in the distribution of fees paid by different transaction types (see Figure 3(d)).
The fact that high-value transactions tend to arrive later in the bidding cycle should have implications for how builders prioritize them within blocks. For Titan, we identify 248 new transactions that arrived after the submission of their first block in the auction cycle, 219 of which were DEX swaps. Titan included the majority of these late-arriving transactions (over 50%) in the top 1.5% percentile positions within the block, confirming the expectation that they were treated as high-priority or high-value. For Rsync, we identify 59 new transactions, 34 of which were DEX swaps. Surprisingly, the median inclusion position for Rsync’s late-arriving transactions is the 68th percentile, that is, the middle of the block. This pattern holds even for Rsync-bot’s unique transactions, which have a median inclusion percentile of 25%.
@@
Of the 669 proposed blocks in our primary dataset, 413 were submitted by Titan Builder and 123 by Rsync Builder. These two builders also account for all of the “high-paying” blocks, as shown in Panel (a) of Figure 4. Panel (b) shows that the revenues from the submitted blocks — measured as the sum of payments to the builder — closely track the builders’ bids over time. The remaining panels of Figure 4 examine the portion of block revenues that Titan and Rsync retain versus those paid to the proposer as bids. These panels show that, as the bidding cycle progresses, both builders gradually reduce the share of revenue they retain, eventually approaching zero. The convergence to zero is faster and more pronounced for Rsync (Panel d) than for Titan (see Panel c).
3.2 Similarity of blocks and transaction sharing
(a) 11 seconds
(b) 12 seconds
(c) 13 seconds
(d) 14 seconds
Figure 5: Proportion of transactions shared between blocks by different builders, at different moments in the bidding cycle.
We begin by analyzing whether blocks submitted by different builders differ in their content. Figure 5 shows the fraction of transactions in each builder’s block that also appear in blocks submitted by other builders. These comparisons are made at various points in the bidding cycle. In this analysis, transactions are identified by their hashes.
We observe that, as the bidding cycle progresses, blocks built by different builders become increasingly dissimilar. This divergence is asymmetric, particularly in the case of Titan. Over time, Titan’s blocks contain fewer transactions that are also present in blocks from other builders, while the reverse is not true: an increasing share of transactions in blocks from other builders are also found in Titan’s blocks. A similar, though weaker, pattern is observed for Rsync. One possible explanation is that losing builders share transactions with builders they expect to win the auction. Equivalently, searchers who initially submit to a builder that appears unlikely to win may later resubmit the same transaction to a builder perceived as more competitive. But it is also possible that some builders, particularly those with integrated searchers, receive more exclusive transactions, which tend to arrive later in the bidding cycle.
To distinguish between these two explanations, we examine whether new transactions received by Rsync or Titan previously appeared in blocks by other builders, where a “new transaction” is a transaction included in a block submitted at least one second after a builder’s initial submission, and not present in that first block. We identify three transactions that appeared only in blocks from Flashbots Builder and Buildernet (a Flashbots-affiliated builder) around 10 seconds into the auction cycle. Then, between 12 and 13 seconds in the bidding cycle, these transactions also appeared in blocks built by Titan. None of them were included on-chain. It is plausible that these transactions were submitted via Flashbots Protect, which only logs transactions that are eventually included, and were later shared with other builders once it became clear that Titan and Rsync were prepared to outbid all non-integrated competitors (we provide further evidence of this behavior in the next section). Overall, while transaction sharing does occur, it is limited in scale and does not fully account for the asymmetries observed in Figure 5.

### v2

@@ 2 Dataset Description and summary statistics
As already mentioned, our primary dataset includes all blocks (winning and non-winning) submitted via a relay called “Agnostic relay” during the study period. [^7] Founded in 2022, Agnostic relay held 20% of the relay market in 2023, declining to around 5% during our study period. We complement our primary dataset with several publicly available data sources:
  • Winning blocks for our study period, including those not supplied by Agnostic relay. Additionally, we track 10 winning blocks beyond our study period to determine whether transactions in our primary dataset were eventually included on-chain.
  • Hashes of blocks submitted to other relays (Flashbots, BloXroute, Manifold, Eden, Ultra Sound, SecureRpc, and Aestus), along with the name of the builders submitting them. Here, a block’s hash serves as its unique identifier, as two blocks that differ in a single transaction have different hashes. We can, therefore, track whether a block in our primary dataset was also submitted to other relays.
  • Data on transactions submitted via the public mempool, obtained from https://mempool.guru/.
  • Data on transactions submitted via two private mempools, MEV Blocker and Flashbots Protect, obtained from https://Dune.com. [^8] Private mempools are services that allow users to send transactions directly to one or more builders without broadcasting them to the public mempool. At present, there are five major private mempool operators. Four of them—MEV Blocker, Flashbots Protect, Merkel, and Blink—state in their documentation that they share transactions privately with multiple builders. Among these, MEV Blocker and Flashbots Protect publicly disclose which transactions included in a winning block were routed through their systems. The fifth operator, Metamask Smart Transaction, does not disclose its sharing policy.
  • Second-by-second price data from Binance (opening price).
We also simulate the non-winning blocks in our dataset to assess how the execution of a transaction — particularly a swap on a blockchain-based financial market — varies depending on the block in which it is included. We simulate the block using an Ethereum Archive Node running the Erigon client, forking at the appropriate block heights.
@@ 2.1 Blocks
Ethereum subdivides time into 12-second slots, during which one block at most can be added to the blockchain. For our purposes, each slot corresponds to a bidding cycle. Slot numbers differ from block numbers; for example, our dataset covers all blocks submitted via the Agnostic relay from slot 10,534,387 to 10,534,425, corresponding to winning blocks 21,322,622 to 21,322,660. For ease of exposition, in what follows, we use winning block numbers whenever possible. Also relevant is that a relay will not hold an auction for a given slot if the proposer assigned to that slot is not registered with the relay. This situation appears in our data: out of the 39 slots we consider, we observe no bids for 11 bidding cycles because the proposers were not connected with the Agnostic relay. [^9] Specifically, we observe no bids for the auction cycles that led to the inclusion of blocks 21,322,624; 21,322,625; 21,322,627; 21,322,629; 21,322,633; 21,322,634; 21,322,636; 21,322,644; 21,322,646; 21,322,647; 21,322,651. For the remaining bidding cycles for which we have data, we find that the blocks submitted to Agnostic relay constitute 28.3% of all blocks submitted to all major relays.
@@
However, when examining the proportion of blocks submitted by the top builders across all relays, we note an imbalance in our primary dataset. Across all relays, Titan Builder, submitted 13,679 blocks (25.6%), Rsync Builder 7,936 blocks (14.9%), Flashbots 2,626 blocks (4.9%), and Beaverbuild 21,447 blocks (40.2%). Hence, Beaverbuild submits proportionally fewer blocks to Agnostic Relay than to other relays. We acknowledge this imbalance as a limitation of our data.
To explore why builders submit fewer blocks to Agnostic Relay than to other relays, we examine the timestamps associated with the blocks that were shared between Agnostic and at least one other relay. We find that only 5% were received by Agnostic before any other relay. More specifically, among the blocks received by both Agnostic and Ultrasound, just 1% of blocks were received by Agnostic first; between blocks received by both Agnostic and BloXroute only 15% were received by Agnostic first. Agnostic, therefore, has a latency disadvantage relative to the other major relays, which may explain why it receives fewer blocks from some builders.
@@ 2.2 Transactions
The first observation is that the set of proposed transactions is very different from the subset of these transactions that appear in winning blocks. The vast majority (87.7%) of the transactions in our primary dataset that appear in winning blocks are users’ transactions, and only a few are searchers’ transactions. [^11] We classify an address as a searcher if it trades on decentralized exchanges (DEXes) using a smart contract whose source code is not disclosed on https://Etherscan.com. In other words, part of the searcher’s logic is embedded in a smart contract but hidden by making only its bytecode available (the bytecode is the compiled, machine-readable version of the smart contract). Swapping on a DEX is necessary for all common types of “searching,” such as sandwich attacks, arbitrage, backrunning, and triggering liquidations on lending protocols. Also, only few (14.4%) of transactions included in winning blocks are swaps. Instead, among the set of proposed transactions, about half are swaps, which are highly concentrated among a few searchers: the top four addresses alone account for 3,653 swap transactions, or approximately 70% of the total swaps.
Users and searchers submit their transactions to builders in different ways. The default option is to send transactions to the Ethereum public mempool, where builders observe them. Doing so, however, exposes users to attacks. [^12] For example, if a user swaps assets on a blockchain-based financial market (an Automated Market Maker, or AMM), a malicious searcher can execute a “sandwich” attack: it front-runs the user’s swap by making the same trade first, then back-runs it with the opposite swap. This allows the attacker to buy low and sell high, forcing the victim to trade at worse prices. To mitigate this risk, users and searchers may instead use private mempools, services that share transactions privately with multiple builders. Furthermore, some searchers may submit their transactions via direct communication with one or more builders. To be included in a block, a transaction should include either a fee or a payment to the “Fee Recipient” address, a block-level variable set by the builder (typically an Ethereum address they control). The sum of the transfers to the “Fee Recipient” address the fees attached to the transactions in a block is the block’s potential revenues, that is, the paiment that a builder receives if the block is successfully added to the blockchain.
@@
A notable feature of our primary dataset is that 2,981 transactions — 35% of all unique transactions — originate from just two addresses. The first address,1313130x68d3A973E7272EB388022a5C6518d9b2a2e66fBf. is responsible for 1,873 transactions, of which only 31 were included in a winning block. Notably, only 6 (out of 1,873) transactions from this address appear in blocks not built by Titan, and all 6 are simple token transfers to Binance. The second address,1414140x51C72848c68a965f66FA7a88855F9f7784502a7F submitted 1,108 transactions, with just 97 included in a winning block. Only 249 (out of 1,108) transactions appear in blocks not built by Rsync, and only 4 appear in blocks not built by either Rsync or Titan. These 4 transactions are also simple token transfers, including two to addresses associated with Wintermute, an algorithmic trading fund that operates Rsync Builder. Based on these patterns, we refer to them as “Titan-bot” and “Rsync-bot”. [^15] Using a different methodology, Heimbach et al. (2024) also identified Rsync-bot as a non-atomic arbitrage bot, while Titan-bot did not yet exist during the period covered by their study. Wu et al. (2025) also identifies Rsync-bot and Titan-bot as non-atomic arbitrage bots “owned by” or “exclusive to” their respective builders. Almost all transactions from these bots involve token swaps on decentralized exchanges (DEXs): 1,864 for Titan-bot and 1,095 for Rsync-bot. We later discuss how these bots often trade on the same DEX pools during the same auction cycle. Based on this activity, we infer that both bots likely engage in non-atomic arbitrage — that is, arbitrage between on-chain and off-chain financial markets — and are often in competition.
@@
To further support our identification of “Rsync-bot” and “Titan-bot,” we analyze whether these searchers submit different transactions to different builders. Using transaction hashes, we find no case in which Rsync-bot sent the same transaction to both Rsync and Titan. However, when comparing transaction logs — i.e., the sequence of operations executed by the transaction and their outputs — we find 203 such cases. The discrepancy between transaction hashes and logs arises because Rsync-bot modifies the fee offered depending on the builder: on average, it pays 18% more to Rsync than to Titan for executing the same transaction. Moreover, we find that transactions sent exclusively to Rsync often include a provision that renders them non-executable if the block’s fee recipient is not Rsync. This behavior further supports our interpretation that Rsync-bot is integrated with Rsync and that the 18% fee differential represents the bot’s targeted profit margin when interacting with non-integrated builders. To perform the same analysis for Titan-bot, we consider winning blocks for our study period that are not in our primary dataset (because they were not submitted via Agnostic relay). Doing so, we find 20 transactions originating from Titan-bot included in blocks won by Beaverbuild. In 3 cases, the same transactions also appear in our dataset in blocks built by Titan. When comparing transaction logs, we find another two transactions in blocks built by Beaverbuild and Titan. In both cases, Titan-bot paid more for inclusion (by 43% and 23%), when the transaction was sent to Titan than Beaverbuild.
@@
Swaps on decentralized exchanges (DEXes) are also the most common type of transaction across all unique transactions in our primary dataset. We identify 5,291 swap transactions, of which 3,919 occur on Uniswap V3, 1,228 on Uniswap V2, 142 on Sushiswap, and 101 on Pancakeswap. [^16] We observe 99 transactions that interact with multiple DEXes, which is why the total number of swaps on individual DEXes exceeds the number of unique swap transactions. Despite their prevalence, only 814 swap transactions — approximately 15% of all swap transactions — were ultimately included on-chain. A small number of addresses, all belonging to searchers, account for the majority of swaps: the top four addresses are responsible for 3,653 swap transactions (about 70%), with Rsync-bot and Titan-bot alone responsible for 2,959 (56%). Swaps are also concentrated among a few pools: the top 11 DEX pools account for 2,872 transactions (54.4%).
@@
We also identify 807 swap transactions that interact with 42 distinct DEX routers, services that optimize swap execution by splitting trades across multiple liquidity pools and/or using intermediate assets. The most frequently used routers are the Uniswap Universal Router (190 transactions), Uniswap V3 Router (172), 1inch V5 Router (52), and Uniswap V2 Router (42). Of these, 594 transactions were included in a winning block, implying that the majority of swaps included on-chain were routed through a DEX router. The remaining 213 router-based transactions were not included on-chain, and the vast majority of these (153) originate from a single user address, 0xf5213a6a2f0890321712520b8048d9886c1a9900. This user submitted 153 transactions calling the Uniswap V3 Router, of which only two were ultimately included in a winning block.
@@ 3 Block 21322649
We now turn to a detailed empirical analysis of a specific auction cycle—namely, the one that resulted in the inclusion of block 21,322,649 on the Ethereum blockchain. This cycle is noteworthy for two reasons. First, bid data across all relays reveals intense competition between Titan Builder and Rsync Builder, with Rsync ultimately winning the auction (see Figures 1). Second, Rsync-bot and Titan-bot actively compete in the same DEX pools, and account for 94%–97% of the total revenue generated by blocks in which they are included. We can therefore closely examine the dynamics of the bidding cycle by analyzing the two bots’ behavior and their respective builders’ bidding strategies.
Figure 1: Builders’ bids across all relays during the auction cycle for Block 21322649
@@ 3.1 Description of the data for the auction cycle of block 21,322,649
For this auction cycle, our primary dataset contains 669 blocks and 504 unique transactions, of which 300 (59%) are swaps. Only 22 swaps (8 from users and 14 from searchers) were included in winning block 21,322,649, which was selected approximately 13 seconds after the previous winning block. An additional 8 swaps were included in winning block 21,322,650, 3 more in winning block 21,322,651 and 1 more in winning block 21,322,652. Notably, all the swaps included in these later blocks are from users. A total of 181 transactions are attributed to Titan-bot and 42 to Rsync-bot, all of which are DEX swaps. None of the Titan-bot transactions were included in a winning block. In contrast, 5 transactions from Rsync-bot were included, all in winning block 21,322,649.
Figure 2: Number of Unique transactions, blocks and swaps vs ETH price change on Binance, during the auction cycle for Block 21322649
@@
As the bidding cycle progresses, we observe a sharp increase in the number of swaps (see Figure 2). This is accompanied by a rise in volume swapped in each block (see Figure 3(a)), as well as in the share of swap volume attributable to Titan-bot and Rsync-bot (see Figure 3(b)). The figure also shows the rate or arrival of different types of transaction, by distinguishing between public transactions (i.e., those that appear in the public mempool), private transactions (i.e., those that do not appear in the public mempool but appear in blocks by multiple builders) and exclusive transactions (i.e., those that do not appear in the public mempool and appear in blocks by a single builder). [^17] Throughout the paper, in the case of public transactions, the relevant timestamp is the first time in which a transaction appears in a proposed block, not the time at which it appears in the public mempool. In contrast to public transactions, which arrive at a relatively uniform rate throughout the bidding cycle, private and, especially, exclusive transactions tend to arrive toward the end of the cycle (see Figure 3(c)). Despite arriving later, exclusive transactions are the most valuable ones, as reflected in the distribution of fees paid by different transaction types (see Figure 3(d)).
@@
Of the 669 proposed blocks in our primary dataset, 413 were submitted by Titan Builder and 123 by Rsync Builder. These two builders also account for all of the “high-paying” blocks, as shown in Panel (a) of Figure 4. Panel (b) shows that the revenues from the submitted blocks — measured as the sum of payments to the builder — closely track the builders’ bids over time. The remaining panels of Figure 4 examine the portion of block revenues that Titan and Rsync retain versus those paid to the proposer as bids. These panels show that, as the bidding cycle progresses, both builders reduce the share of revenue they retain, eventually approaching zero. The convergence to zero is faster and more pronounced for Rsync (Panel d) than for Titan (see Panel c).